One state lottery has 1100 prizes of $1, 120 prizes of $10, 20 prizes of $55, 5 prizes of $275, 2 prizes of $1300, and 1 prize of $2700. Assume that 35000 lottery tickets are issued and sold for $1 1. What is the lottery's expected profit per ticket? 2. What is the lottery's standard deviation of profit per ticket?
Answer #1
x P(x) xP(x) x2P(x)
0 0.0314 0.0000 0.0000
9 0.0034 0.0309 0.2777
54 0.0006 0.0309 1.6663
274 0.0001 0.0391 10.7251
1299 0.0001 0.0742 96.4229
2699 0.0000 0.0771 208.1315
-1 0.9643 -0.9643 0.9643
total -0.7121 318.1879
E(x) =μ= ΣxP(x) = -0.7121
E(x2) = Σx2P(x) = 318.1879
Var(x)=σ2 = E(x2)-(E(x))2= 317.6807
std deviation=         σ= √σ2 = 17.8236

1) lottery's expected profit per ticket = -0.7121

2) lottery's standard deviation of profit per ticket =17.8236